if the mass of the earth is twice as much as at present then the time period of moon revolving at the same distance at present will be...

Grade:12th pass

1 Answers

Rama sai rohan
6 years ago
As the moon is revolving arround earth So centripetal force =gravitational force "M"is mass of eart "m" is mass of moonR is the distance between them(mv^2)/R=GMm/R^2GM/R=V^2 here V=2πR/TGM/R =4π^2R^2/T^2 T^2=(4π^2R^3)/(GM)T is proportional to 1/√MAs" M" is doubled time will be 1/√2 times of initial